# Can Newtonian Gravity Produce Quantum Entanglement?

### v2

@@ IV Entangling quantum bodies by Newtonian gravity ?
$U_{T}:={\cal T}e^{-i\int_{-\infty}^{\infty}w_{T}(t)\hat{H}_{QQ}dt}\;,$ (17)
where $\hat{H}_{QQ}$ can be either $H_{QQ}^{\rm mini}$, $H_{QQ}^{\rm semi}$ or $H_{QQ}^{\rm stoc}$, and $w_{T}$ is some window function, which is chosen to be $w_{T}(t)=e^{-(t/T)^{2}}/(\pi/2)^{1/4}$ such that $\int_{-\infty}^{\infty}|w_{T}(t)|^{2}dt=T$. The initial state of the two quantum bodies is just
@@
with $\theta^{(a)}\neq n\pi/2$ for $n\in\mathbb{Z}$. Otherwise, the quantum bodies will be in their eigenstates, acting as classical bodies.
@@ IV.1 Entanglement production in mini-superspace approach
For the mini-superspace approach with $\hat{H}_{QQ}^{\rm mini}$ of (13), then
@@
$\lambda:=-H_{QQ}^{c}\int_{-\infty}^{\infty}dt\;w_{T}(t)=-(2\pi)^{1/4}H_{QQ}^{c}T\sim{\cal O}(1){G\delta M_{1}\delta M_{2}\over r_{12}}\Big({b\over r_{12}}\Big)^{4}T\;.$ (20)
For simplicity, we consider the following initial state with $\theta^{(1)}=\theta^{(2)}={\pi\over 4}$ and $\phi^{(1)}=\phi^{(2)}=0$, i.e.,
$|{\rm init}\rangle=|++\rangle:=|+;1\rangle\otimes|+;2\rangle\;,\qquad{\rm with}\quad|+;a\rangle:=|\theta^{(a)}=\pi/4,\phi^{(a)}=0;Q^{(a)}_{ij}\rangle\;$. (21)
Further define and note
$|--\rangle:=|-;1\rangle\otimes|-;2\rangle\qquad{\rm with}\qquad|-;a\rangle:=\hat{Z}^{(a)}|+;a\rangle\quad{\rm and}\quad\langle\pm;a|\mp;a\rangle=0\;.$ (22)
Similarly, one can define $|+$-$\rangle$ and $|-+\rangle$. Then,
@@
The Schmidt rank generally increases by one after evolution unless $\lambda=n\pi/2$ with $n\in{\mathbb{Z}}$, so that the mini-superspace quantum gravity entangles the quantum bodies. This is consistent with the GIE proposal, i.e., the production of quantum entanglement can be used as a probe to quantum gravity.
@@ IV.2 No entanglement production in semiclassical and stochastic gravity approaches
with $\Theta^{(a)}=\theta^{(a)}$ for semiclassical gravity, and $\tilde{\theta}^{(a)}$ for stochastic gravity. The product of local unitaries cannot produce entanglement as can be seen as follows,
$|{\rm final}\rangle=e^{-\lambda_{2}\hat{Z}^{(1)}}|\theta^{(1)},\phi^{(1)};Q^{(1)}_{ij}\rangle\otimes e^{-\lambda_{1}\hat{Z}^{(2)}}|\theta^{(2)},\phi^{(2)};Q^{(2)}_{ij}\rangle\;$, (26)
@@ IV.3 Confusing GIE-violating result due to perturbative calculations
The recent paper Aziz and Howl (2025) claims that classical gravity can produce quantum entanglement, which is based on perturbative calculations in the context of the stochastic gravity approach. We now demonstrate a subtlety for reaching this GIE-violating result. For simplicity, we consider the initial state of (21) so that $\lambda_{a}$’s of (25) vanish as $\cos 2\theta^{(a)}\big|_{\theta^{(a)}=\pi/4}=0$ for semiclassical gravity approach, but not for stochastic gravity approach due to nontrivial $\delta\theta$. So, there is no evolution for the semiclassical gravity in this case. On the other hand, for the stochastic gravity, we have
@@
Now, we can explain the possible origin of claiming the GIE-violating result as follows. In the perturbative approach as performed in Aziz and Howl (2025), the resultant final state is obtained up to some order of $G_{N}$. This corresponds to expanding the right side of (27) in order of $\lambda_{a}$ as $\lambda_{a}\sim{\cal O}(G_{N})$. Suppose we are interested in the final state up to ${\cal O}(G_{N})$, then we have
$|{\rm final}\rangle=\big(1-i\lambda_{2}\hat{Z}^{(1)}+\cdots\big)|+;1\rangle\otimes\big(1-i\lambda_{2}\hat{Z}^{(2)}+\cdots\big)|+;2\rangle\simeq|++\rangle-i\lambda_{1}|+-\rangle-i\lambda_{2}|-+\rangle+{\cal O}(G_{N}^{2})\;.$ (28)
@@
Acknowledgements. The work of FLL is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. The work of SM is supported by ANID FONDECYT Postdoctorado grant number 3240055.

### v3

@@ IV Does Newtonian gravity entangle quantum bodies?
$U_{T}:={\cal T}\exp\left({-i\int_{-\infty}^{\infty}w_{T}(t)\hat{H}_{QQ}dt}\right)\;,$ (15)
where $\hat{H}_{QQ}$ corresponds to either the mini-superspace ($H_{QQ}^{\rm mini}$), semiclassical ($H_{QQ}^{\rm semi}$), or stochastic ($H_{QQ}^{\rm stoc}$) Hamiltonian, and $\cal T$} is the time-odering operator. The window function is chosen as $w_{T}(t)=e^{-(t/T)^{2}}(\pi/2)^{-1/4}$, satisfying the normalization $\int|w_{T}|^{2}dt=T$. We choose the initial state of the two quantum bodies to be a product state
@@
requiring $\theta^{(a)}\neq n\pi/2$ ($n\in\mathbb{Z}$) to ensure the bodies are in macroscopic superpositions rather than eigenstates of the interaction basis, otherwise it would mimic classical behavior.
@@ IV.1 Entanglement production in mini-superspace approach
For the mini-superspace approach with $\hat{H}_{QQ}^{\rm mini}$ is defined in (11), then the time evolution operators simplifies to
@@
$\lambda$ $:=$ $-H_{QQ}^{c}\int_{-\infty}^{\infty}dt\;w_{T}(t)\;=-(2\pi)^{1/4}H_{QQ}^{c}T\;,$ (18)
$\sim$ ${\cal O}(1){G\delta M_{1}\delta M_{2}\over r_{12}}\Big({b\over r_{12}}\Big)^{4}T\;.$ (19)
Considering the specific initial configuration $\theta^{(a)}=\pi/4$ and $\phi^{(a)}=0$, denoted as $|{\rm init}\rangle=|++\rangle\equiv|+;1\rangle\otimes|+;2\rangle$, we define the orthogonal states $|-;a\rangle:=\hat{Z}^{(a)}|+;a\rangle$. The evolved state becomes
$|{\rm init}\rangle=|++\rangle:=|+;1\rangle\otimes|+;2\rangle\;,$ (20)
where
$|+;a\rangle:=|\theta^{(a)}=\pi/4,\phi^{(a)}=0;Q^{(a)}_{ij}\rangle\;$. (21)
Furthermore, we define
$|--\rangle:=|-;1\rangle\otimes|-;2\rangle\;,$ (22)
where
$|-;a\rangle:=\hat{Z}^{(a)}|+;a\rangle\quad{\rm and}\quad\langle\pm;a|\mp;a\rangle=0\;.$ (23)
for $a=1,~\text{and}~2$. Similarly, one can define $|+$-$\rangle$ and $|-+\rangle$. Then, the time-evolved final states becomes
@@
For $\lambda\neq n\pi/2$ where $n\in\mathbb{Z}$, the Schmidt rank increases to two, confirming that the mini-superspace formulation—representing quantized gravity—entangles the bodies. This aligns with the GIE protocol, establishing entanglement generation as a probe for quantum gravity.
@@ IV.2 No entanglement production in semiclassical and stochastic gravity approaches
with $\Theta^{(a)}=\theta^{(a)}$ for semiclassical gravity, and $\tilde{\theta}^{(a)}$ for stochastic gravity. The product of local unitaries cannot produce entanglement, as can be seen as follows,
$|{\rm final}\rangle=$ $e^{-\lambda_{2}\hat{Z}^{(1)}}|\theta^{(1)},\phi^{(1)};Q^{(1)}_{ij}\rangle$ (27)
$\otimes\;e^{-\lambda_{1}\hat{Z}^{(2)}}|\theta^{(2)},\phi^{(2)};Q^{(2)}_{ij}\rangle\;$,
@@ IV.3 Confusing GIE-violating result due to perturbative calculations
The recent paper [Aziz:2025ypo] claims that classical gravity can produce quantum entanglement, which is based on perturbative calculations in the context of the stochastic gravity approach. We now demonstrate a subtlety for reaching this GIE-violating result. For simplicity, we consider the initial state of (20) so that $\lambda_{a}$’s of (26) vanish as $\cos 2\theta^{(a)}\big|_{\theta^{(a)}=\pi/4}=0$ for semiclassical gravity approach, but not for stochastic gravity approach due to nontrivial $\delta\theta$. So, there is no evolution for the semiclassical gravity in this case. On the other hand, for the stochastic gravity, we have
@@
Now, we can explain the possible origin of claiming the GIE-violating result as follows. In the perturbative approach as performed in [Aziz:2025ypo], the resultant final state is obtained up to some order of $G_{N}$. This corresponds to expanding the right side of (28) in order of $\lambda_{a}$ as $\lambda_{a}\sim{\cal O}(G_{N})$. Suppose we are interested in the final state up to ${\cal O}(G_{N})$, then we have
$|{\rm final}\rangle$ $=$ $\big(1-i\lambda_{2}\hat{Z}^{(1)}+\cdots\big)|+;1\rangle\otimes\big(1-i\lambda_{2}\hat{Z}^{(2)}+\cdots\big)|+;2\rangle\;,$ (29)
$\simeq$ $|++\rangle-i\lambda_{1}|+-\rangle-i\lambda_{2}|-+\rangle+{\cal O}(G_{N}^{2})\;.$
@@
V Discussion and conclusion
In this article, we systematically investigated the capacity of Newtonian gravity to induce quantum entanglement between mesoscopic quantum bodies, specifically modeled as mass quadrupoles. Our theoretical framework analyzed three distinct interaction scenarios to elucidate the mechanisms of entanglement generation under gravitational influence: (1) mini-superspace approach, (2) semiclassical gravity, and (3) stochastic gravity.
Our results provide robust support for the gravity-induced entanglement (GIE) hypothesis, reinforcing the postulate that classical gravity cannot mediate quantum entanglement. We demonstrated that the mini-superspace approach successfully facilitates entanglement production. This is achieved by quantizing the parity of gravitational tidal fields sourced by the mass quadrupoles. Conversely, both semiclassical and stochastic gravity models failed to generate entanglement, yielding only final product states. These findings confirm the necessity of a quantum gravitational framework for the effective production of entanglement, aligning with the core implications of the GIE protocol.
A critical aspect of our work involves addressing recent assertions that classical gravity might induce entanglement under specific conditions. We clarified that such claims likely arise from perturbative calculations that neglect higher-order cross-terms. By isolating these artifacts, we demonstrated how truncations in expansions can lead to misleading interpretations, thereby resolving potential confusion regarding the quantum nature of gravitational interactions.
Our results establish a foundation for further exploration into the interplay between gravity and quantum systems. As experimental techniques continue to advance, developing concrete models grounded in rigorous theoretical frameworks will be essential to probing the intricate link between gravity and quantum mechanics, ultimately deepening our understanding of fundamental physics.
VI Acknowledgements
SM is grateful to Ignacio Araya and Giorgos Anastasiou for their hospitality at Universidad Andrés Bello and Universidad Adolfo Ibañez, respectively. The work of FLL is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. The work of SM is supported by ANID FONDECYT Postdoctorado grant number 3240055.